OpenStudy (anonymous):

Evaluate the triple integral (1+z^2)dV over D, where D is the solid region bounded below by the upper portion of the cone z^2=3x^2+3y^2 and above by the sphere x^2+y^2+z^2=4

OpenStudy (anonymous):

z would go from \[\int _{\sqrt{3 x^2+3 y^2}}^{-\sqrt{-x^2-y^2+4}}\]

OpenStudy (anonymous):

so should i integrate z with those limits and then change to polar coordinates?